OpenStudy (anonymous):

Not sure if I already posted this If q = 1/2 when r = 1/3 and q varies inversely with r, then find q when r = 6. A. 1/6 B. 6 C. 1/36 D. 36

OpenStudy (anonymous):

(q)(r) = k -> (1/2)(1/3) = k -> k = 1/6 (q)(r) = 1/6 -> (q)(6) = 1/6 -> q = (1/6) / 6 q = 1/36

OpenStudy (anonymous):

It's the "inverse proportionality" situation again. When q gets bigger, then r gets smaller.

OpenStudy (anonymous):

alright thanks :D